OpenStudy (anonymous):

Sixteen squared divided by four to the fourth power, write as a given expression and then evaluate.

OpenStudy (raden):

it means : 16^2 / 4^4 = (2^4)^2 / (2^2)^4 use the property of exponent that : (a^m)^n = a^(m*n)

OpenStudy (raden):

so, it can be : 16^2 / 4^4 = (2^4)^2 / (2^2)^4 = 2^8/2^8 i think u will get its answer, now